### v4

# Fair Column Subset Selection

Abstract
The problem of column subset selection asks for a subset of columns from an input matrix such that the matrix can be reconstructed as accurately as possible within the span of the selected columns. A natural extension is to consider a setting where the matrix rows are partitioned into two groups, and the goal is to choose a subset of columns that minimizes the maximum reconstruction error of both groups, relative to their respective best rank-$k$ approximation. Extending the known results of column subset selection to this fair setting is not straightforward: in certain scenarios it is unavoidable to choose columns separately for each group, resulting in double the expected column count. We propose a deterministic leverage-score sampling strategy for the fair setting and show that sampling a column subset of minimum size becomes NP-hard in the presence of two groups. Despite these negative results, we give an approximation algorithm that guarantees a solution within 1.5 times the optimal solution size. We also present practical heuristic algorithms based on rank-revealing QR factorization. Finally, we validate our methods through an extensive set of experiments using real-world data.
1 Introduction
Dimensionality reduction techniques such as principal component analysis (PCA) and non-negative matrix factorization have proven useful for machine learning and data analysis tasks (Pearson, 1901; Lee and Seung, 1999). Such tasks include feature selection, feature extraction, noise removal and data visualization, among others. These techniques are also commonly employed as components of larger machine learning (ML) pipelines to simplify data through lower-dimensional representations. However, when the data is divided into subsets (or groups), an inaccurate low-dimensional representation of any subset can perpetuate inaccuracies in subsequent downstream tasks. Therefore, there is increasing emphasis on developing techniques that produce accurate representations for all the different groups.
Notably, Samadi et al. (2018) showed that a well-known dimensionality reduction technique, PCA, may incur higher average reconstruction error for a subset of the population than the rest, even when the groups are of similar size. They proposed fair variants of PCA, where the objective is to minimize the maximum reconstruction error for any group. However, a drawback of PCA is that the results are often hard to interpret, since its output consists of abstract attributes that might not necessarily be part of the input data. As an alternative to PCA, one may ask for algorithms that choose a (small) subset of the original attributes of the dataset to act as a low dimensional representation.
In the column subset selection problem (CSS), we are given a data matrix and seek a representative subset of its columns. The quality of the solution is measured by the residual norm when the input matrix is projected onto the subspace spanned by the chosen columns. CSS has been extensively studied, and polynomial-time approximation algorithms are known for different quality criteria (Deshpande et al., 2006; Deshpande and Rademacher, 2010; Boutsidis et al., 2009, 2014; Papailiopoulos et al., 2014; Altschuler et al., 2016).
In this work, we study the fair column subset selection problem (FairCSS), where the rows of the data matrix are partitioned into two groups. The goal is to choose a subset of columns that minimize the maximum reconstruction error of both groups, through a min-max objective. This distinguishes our problem from conventional CSS methods that aim to minimize the reconstruction error of the entire data matrix and may neglect either group. Here, we focus on the two-group setting. Generalizing our results to an arbitrary number of groups is a non-trivial technical leap and thus left as future work. Nevertheless, binary protected attributes are encountered often in practice and are commonly the focus of works on algorithmic fairness (Samadi et al., 2018; Chierichetti et al., 2017).
To further motivate the study of FairCSS, consider its application in drug discovery. Suppose we are given a large dataset of medical records, where each row represents a patient and the columns contain genetic information indicating presence (absence) of a gene expression with binary value $1$ (00). Due to the size of the genetic data, it is reasonable to ask for a succint representation of the dataset. Using PCA would yield an arbitrary subspace of the data that is hard to interpret. On the other hand, classical CSS would return a representative set of columns (genes), however, if there is bias in the medical records then the reconstruction error could skew favouring a (majority) group. Thus, any subsequent insights derived from the biased (succint) data could also perpetuate these biases. It is possible that, a drug discovered using this biased data is more (less) effective for men (women), based on the data majority. Our proposed method, however, returns a subset of columns that are representative of both groups, thus mitigating potential biases and ensuring fair representation in subsequent analysis.
The two-group setting of CSS introduces significant challenges, making it necessary to view the problem through a new lens. In CSS, finding the optimal $k$-column subset is NP-hard (Shitov, 2021) and thus, polynomial-time approximation algorithms are usually sought. A factor of $O(k)$ with respect to the best rank-$k$ approximation is possible if we are allowed to choose at most $k$ columns (Deshpande and Rademacher, 2010), and better results can be obtained if we allow $c\geq k$ columns (Boutsidis et al., 2008). In the two-group fair setting, similar approximation bounds can be reproduced if we allow twice as many columns in the solution; by optimizing for each group separately. However, addressing the two groups separately can raise ethical and legal concerns (Lipton et al., 2018; Samadi et al., 2018), potentially conflicting with the principles of demographic parity and equal treatment (Barocas et al., 2019). Further, this approach may inadvertently contribute to segregation and stereotyping. Therefore, our focus is to find a common subset of representative columns for both groups.
Unfortunately, as we show with an example (see § 3), in certain scenarios it may not be possible to do better than choosing twice as many columns. Can we hope to obtain guarantees of any kind in this fair setting? We answer this question affirmatively. To achieve this, we adapt leverage score sampling to our fair setting. Leverage scores are obtained using the SVD of the matrix, and they are used to find provably good column subsets for CSS. In particular, Papailiopoulos et al. (2014) showed that by choosing columns with highest leverage scores such that their sum adds to a predefined threshold $\theta$, a relative-error approximation is possible. In standard CSS, finding a column subset of minimum size satisfying threshold $\theta$ is trivial, by sorting the leverage scores. In the fair setting however, finding a subset of leverage scores of minimum size satisfying threshold $\theta$ for both groups is NP-hard. While the original result can be extended to the fair setting by solving for both groups independently and doubling the column count, we present an efficient algorithm that achieves this with essentially $1.5$ as many columns. Whether this factor can be improved is left as an open question. Finally, we introduce efficient heuristics for the problem based on QR factorizations with pivoting, and assess their performance in our empirical evaluation. Our contributions are summarized as follows:
  • $\bullet$
    We introduce the novel problem of fair column subset selection, where two groups are present in the data and the maximum reconstruction error of the two groups must be minimized.
  • $\bullet$
    We extend the approach of deterministic leverage score sampling to the two-group fair setting. We show that the smallest column subset that achieves the desired guarantees is NP-hard to find, and give a polynomial-time algorithm wih relative-error guarantees with a column subset of essentially $1.5$ times the minimum possible size.
  • $\bullet$
    We present efficient heuristics based on QR factorizations with column pivoting.
  • $\bullet$
    We empirically evaluate our algorithms on real-world data and show they are able to select fair columns with high accuracy. Further, we analyse the price of fairness for our formulation.
The paper is organised as follows: Section 2 covers related work. Section 3 outlines our problem. We adapt deterministic leverage score sampling to the fair setting in Section 4. Algorithms based on QR decomposition are given in Section 5. Experimental results are presented in Section 6. Finally, we conclude in Section 7.
2 Related Work
Our work builds upon related work in the areas of algorithmic fairness and column subset selection.
Algorithmic fairness. The influential work of Dwork et al. (2012) established a formal notion of fairness in algorithmic decision-making, which has served as a foundation for subsequent research in the field. Pedreschi et al. (2008) addressed discrimination in data mining, while Kamiran and Calders (2010, 2012) proposed a framework for mitigating discrimination in classification. Since then, there has been extensive research on algorithmic fairness from various disciplines, including economics, game theory, statistics, ethics, and computer science (Rambachan et al., 2020; Sætra et al., 2022; Raji and Buolamwini, 2022; Buolamwini and Gebru, 2018). Fair objectives have been introduced into many classical computer science problems (Chierichetti et al., 2017; Samadi et al., 2018; Anagnostopoulos et al., 2020; Ghadiri et al., 2021; Froese et al., 2022; Zemel et al., 2013; Zafar et al., 2017; Zehlike et al., 2017). Fair clustering share similarities to fair CSS, with a key distinction: the former asks for a subset of fair representative data points (rows), while the later focuses on finding a fair representative column subset (attributes) Chierichetti et al. (2017); Chhabra et al. (2021); Ghadiri et al. (2021). For further reading see Pessach and Shmueli (2022); Mitchell et al. (2021); Kleinberg et al. (2018).
Column Subset Selection. CSS can be viewed as a method for feature selection. Early work in CSS traces back to the numerical linear algebra community, and the seminal works of Golub (1965) on pivoted QR factorizations, which was followed by works by Chan (1987); Hong and Pan (1992); Chan and Hansen (1994) addressing the problem of finding efficient rank revealing QR factorizations (RRQR). Recently, CSS has attracted interest in the computer science community, with approaches that combine sampling with RRQR (Boutsidis et al., 2009; Papailiopoulos et al., 2014), while greedy methods have proved highly effective (Altschuler et al., 2016; Farahat et al., 2015).
Fair PCA. Closest to our work is fair principal component analysis (FairPCA) by Samadi et al. (2018), where the goal is to find a low-dimensional projection of the data to optimize a min-max objective that ensures fairness with respect to two groups. Tantipongpipat et al. (2019) extended FairPCA to a multi-objective setting. Olfat and Aswani (2019) presented a polynomial-time algorithm using convex programming for the general case with multiple groups. Similar to FairPCA, we employ a min-max objective, but in contrast, we want to find a subset of actual columns to approximate the reconstruction error of both groups. In fact, we show that the problem we define, fair CSS, is NP-hard for any number of groups (see § 3).
3 Problem statement
In this section, we describe relevant terminology and formally define CSS before introducing fair CSS. For a positive integer $n$ we denote $[n]=\{1,\ldots,n\}$. For any matrix $C$, $P_{C}$ denotes the projection operator onto the subspace spanned by the columns of $C$.
Problem 1 (Column Subset Selection (CSS)).
Given a matrix $M\in\mathbb{R}^{m\times n}$ and a positive integer $k$. The goal is to choose $k$ columns of $M$ to form a matrix $C\in\mathbb{R}^{m\times k}$ such that the reconstruction error
$loss(M,C)=\|M-P_{C}M\|_{F}$
is minimized, where $\|\cdot\|_{F}$ denotes the Frobenius norm.
Fairness. A solution to CSS gives the best possible column approximation for the matrix $M$ overall. However, if the matrix rows are partitioned into two groups, it is possible that the reconstruction error shows significant disparity when measured on each group separately, for instance, if one of the groups is a minority. In the fair setting, our goal is to choose a subset of columns that achieves good reconstruction error for both groups. To formalize this, assume that the rows of matrix $M$ are partitioned into two subsets $A$ and $B$. Subscripting a matrix by $A$ or $B$ denotes the rows corresponding to $A$ or $B$. For ease of notation, we override $M_{A}=A$ and $M_{B}=B$.
Before presenting our fair CSS formulation, we make two considerations. In CSS, the optimal projection is obtained as $P_{C}=CC^{+}$, where $C^{+}$ is the pseudoinverse of $C$. In our setting, even though we choose a common column subset for both groups, the optimal reconstruction error for each group is attained by a different projection operator, that is, $C_{A}C_{A}^{+}A$ and $C_{B}C_{B}^{+}B$. Second, we are interested in minimising the relative loss of each group, with respect to the optimal reconstruction error, which can be obtained using the best rank-$k$ approximation. This is to avoid excessively penalizing either of the two groups, when the other group does not have a good low-rank representation. Note that measuring the reconstruction error relatively to the best rank-$k$ approximation is common in CSS literature. See Boutsidis et al. (2008); Papailiopoulos et al. (2014).
Definition 1 (Relative group-wise reconstruction error).
Given a matrix $M\in\mathbb{R}^{m\times n}$, a row subset $A\in\mathbb{R}^{a\times n}$, with $A_{k}$ its best rank-$k$ approximation, and a matrix $C\in\mathbb{R}^{m\times k}$ formed by choosing $k$ columns of $M$. The relative reconstruction error of $A$ is:
$Nloss_{A}(M,C)=\frac{\|A-C_{A}C_{A}^{+}A\|_{F}}{\|A-A_{k}\|_{F}}.$
Now we formally define our problem.
Problem 2 (FairCSS-MinMax).
Given a matrix $M\in\mathbb{R}^{m\times n}$ with row partition $A,B$, and a positive integer $k$. The goal is to choose $k$ columns of $M$ to form a matrix $C\in\mathbb{R}^{m\times k}$ that optimizes the objective:
$\min_{\begin{subarray}{c}C\in\mathbb{R}^{m\times k}\\ C\subset M\end{subarray}}\max\biggl{\{}\frac{\|A-C_{A}C_{A}^{+}A\|_{F}}{\|A-A_{k}\|_{F}},\frac{\|B-C_{B}C_{B}^{+}B\|_{F}}{\|B-B_{k}\|_{F}}\biggl{\}}.$
Thus, we search for a subset of columns that minimizes the maximum error of either groups. Observe that when groups $A$ and $B$ are identical, FairCSS-MinMax is equivalent to CSS, which is known to be NP-hard (Shitov, 2021)[Theoreom 2.2]. The hardness results extend to FairCSS-MinMax. In line with preceding arguments, FairCSS-MinMax is NP-hard for any number of groups.
Observation 1.
FairCSS-MinMax is NP-hard.
Limitations. We show that, in some scenarios it is not possible to do better than solving for two groups separately. In particular, any algorithm that attempts to solve CSS for more than one group, i.e., to achieve errors smaller than $\|A\|_{F}^{2}$ and $\|B\|_{F}^{2}$ for $A$ and $B$, respectively, cannot achieve meaningful bounds on the relative-error with respect to both $\|(A)_{k}\|_{F}^{2}$ and $\|(B)_{k}\|_{F}^{2}$. Consider the following example:
$\left(\begin{matrix}X_{A}&0_{A}\\ 0_{B}&X_{B}\end{matrix}\right),$
where $X_{A},X_{B}$ are matrices of rank $n>k$. If we pick less than $2k$ columns, the error is at least $\min\left\{\frac{\sum_{i=k}^{n}\sigma_{i}^{2}(A)}{\sum_{i=k+1}^{n}\sigma_{i}^{2}(A)},\frac{\sum_{i=k}^{n}\sigma_{i}^{2}(B)}{\sum_{i=k+1}^{n}\sigma_{i}^{2}(B)}\right\}$, where $\{\sigma_{1},\dots,\sigma_{n}\}$ are the singular values with $\sigma_{i}\geq\sigma_{i+1}$. Thus, the relative error can be unbounded if the rank of either submatrix is numerically close to $k$, i.e., if $\sigma_{k}(A)\gg\sigma_{k+1}(A)$ or $\sigma_{k}(B)\gg\sigma_{k+1}(B)$. Clearly, the only way to prevent this is to pick $2k$ columns. A similar result is yielded by matrices where the blocks of zeroes are replaced by small values. Despite this, in the following section we present an algorithm with a bounded-error relative to the best rank-$k$ approximation, by relaxing the requirement on the number of selected columns.
4 Pairs of leverage scores
In this section, we discuss leverage scores in the context of the two-group fair setting. Leverage scores are a pivotal concept that is extensively studied in CSS literature, as they provide valuable insights for selecting column subsets with provable approximation guarantees. We begin by discussing the results of Papailiopoulos et al. (2014) on leverage score sampling. Next, we show that sampling a subset of columns of minimum size in the two-group fair setting is NP-hard. Finally, we present an approximation algorithm that samples $1.5$ times the size of an optimal solution. An useful concept that is extensively studied in CSS literature is leverage scores, which is defined as follows:
Definition 2 (Leverage scores).
Let $V^{(k)}\in\mathbb{R}^{n\times k}$ denote the top-$k$ right singular vectors of a $m\times n$ matrix $M$ with rank $\rho=rank(M)\geq k$. Then, the rank-$k$ leverage score of the $i$-th column of $M$ is defined as:
[…]
Here, $[V^{(k)}]_{i,:}$ denotes the $i$-th row of $V^{(k)}$.
Leverage scores are used to find a solution with approximation guarantees for CSS. In particular, we focus on the following result by Papailiopoulos et al. (2014).
Theorem 1 (Papailiopoulos et al. (2014)).
Given a matrix $M\in\mathbb{R}^{m\times n}$ and an integer $k<\text{rank}(M)$. Let $\theta=k-\epsilon$ for some $\epsilon\in(0,1)$ and $S$ be a subset of column indices such that $\sum_{i\in S}\ell_{i}^{(k)}\geq\theta$, and $C\in\mathbb{R}^{m\times k}$ be the matrix of $M$ formed by choosing the columns with indices in $S$. Then we have that
$\|M-CC^{+}M\|_{F}^{2}\leq(1-\epsilon)^{-1}\|M-M_{k}\|_{F}^{2}.$
[…]
Depending on the leverage score distribution, sometimes it may be necessary to pick more than $k$ columns to satisfy the threshold $\theta$, and $\Omega(n)$ in the worst case. Nevertheless, when the leverage scores follow power-law decay, a small factor of $k$ suffices (Papailiopoulos et al., 2014)[Theorem 3].
Fair deterministic leverage-score sampling. In order to achieve approximation guarantees for both groups, the leverage scores of both $A$ and $B$ must sum to at least $\theta$, individually. While seeking the minimum number of columns to satisfy the threshold is trivial in the single-group setting, it becomes NP-hard in the presence of two groups. Let us formally define the problem and analyse its complexity.
Problem 3 (Min-FairnessScores).
Given matrices $A\in\mathbb{R}^{m_{A}\times n}$ and $B\in\mathbb{R}^{m_{B}\times n}$, $k\in\mathbb{N}:0<k<\text{rank}(A),\text{rank}(B)$ and a threshold $\theta=k-\epsilon$ for some $\epsilon\in(0,1)$, find the smallest set of indices $S\subseteq[n]$ such that:
$\sum_{i\in S}\ell_{i}^{(k)}(A)\geq\theta\text{~{}and~{}}\sum_{i\in S}\ell_{i}^{(k)}(B)\geq\theta.$
If we find a subset of columns that satisfy both inequalities above, then due to Theorem 1, we have that,
$\|A-C_{A}C_{A}^{+}A\|_{F}\leq(1-\epsilon)^{-1/2}\|A-A_{k}\|_{F}\implies Nloss_{A}(M,C)\leq(1-\epsilon)^{-1/2}.$
and similarly, $Nloss_{B}(M,C)\leq(1-\epsilon)^{-1/2}.$
A solution to Problem 3 gives us an upper bound on the reconstruction error. Unfortunately, MinFairnessScores is NP-hard.
Theorem 2.
MinFairnessScores is NP-hard.
Proof.
To establish hardness we reduce the equal cardinality partition problem (EqCardPartition) to a decision version of MinFairnessScores, called $c$-FairnessScores. The decision version asks to find exactly $c$ indices i.e., $S\subseteq[n]$, $|S|=c$ such that $\sum_{i\in S}\ell_{i}^{(k)}(A)\geq\theta$ and $\sum_{i\in S}\ell_{i}^{(k)}(B)\geq\theta$. Given a set $Z=\{p_{1},\ldots,p_{n}\}$ of $n$ positive integers, EqCardPartition asks to partition $Z$ into two disjoint subsets $X,Y$ such that $X\cup Y=Z$, $|X|=|Y|$ and $\sum_{p_{i}\in X}p_{i}=\sum_{p_{j}\in Y}p_{j}$. EqCardPartition is known to be NP-complete (Garey and Johnson, 1979, SP12).
Given an instance of EqCardPartition $(Z,n)$, we reduce it to a $c$-FairnessScores instance $(A,B,\theta,c)$ as follows. Let $s=\sum_{p_{i}\in Z}p_{i}$ and $M\gg s$ be some constant. Let $A=[\sqrt{p_{1}},\ldots,\sqrt{p_{n}}]$ and $B=\big{[}\sqrt{M-\frac{p_{1}}{s}},\ldots,\sqrt{M-\frac{p_{n}}{s}}\big{]}$ be input matrices such that $A\in\mathbb{R}^{n\times 1}$, $B\in\mathbb{R}^{n\times 1}$. Finally, we set $\theta=1/2$ and $c=n/2$. We claim that EqCardPartition is a yes instance if and only if $c$-FairnessScores is a yes instance. The reduction is polynomial in the input size.
To make the reduction work, we need a way to map the values in the instance EqCardPartition to leverage scores, which are obtained from SVD. The most straightforward way is to compute the rank-$1$ leverage scores. For all $i\in[n]$ we can get,
$\big{(}\ell_{i}^{(1)}(A),\ell_{i}^{(1)}(B)\big{)}=\Big{(}\frac{p_{i}}{s},\frac{M-\frac{p_{i}}{s}}{nM-1}\Big{)}.$
For ease of notation, let $\alpha_{i}=\ell_{i}^{(k)}(A)$ and $\beta_{i}=\ell_{i}^{(k)}(B)$.
Let $c$-FairnessScores be a yes instance. Then we have a subset $S\subseteq[n]$ such that $|S|=n/2$,
$\sum_{i\in S}\alpha_{i}=\frac{\sum_{i\in S}p_{i}}{s}\geq\frac{1}{2}\text{~{}which implies }\sum_{i\in S}p_{i}\geq\frac{s}{2},$
$\sum_{i\in S}\beta_{i}=\frac{\sum_{i\in S}(M-\frac{p_{i}}{s})}{nM-1}\geq\frac{1}{2}\text{~{}which implies }\sum_{i\in S}p_{i}\leq\frac{s}{2}.$
For both of the above inequalities to hold simultaneously, we must have $\sum_{i\in S}p_{i}=s/2$. Thus, $X=\{p_{i}:i\in S\},Y=\{p_{i}:i\in[n]\setminus S\}$ is a solution to EqCardPartition.
For the sake of contradiction, assume that $c$-FairnessScores is a no instance and $X,Y$ are a solution to EqCardPartition. Let $S$ be the set of indices of elements in $X$. We can choose $S$ as solution for $c$-FairnessScores since $|S|=n/2$,
$\sum_{i\in S}\alpha_{i}=\sum_{i\in S}p_{i}/s=\frac{1}{2},\text{ and}\sum_{i\in S}\beta_{i}=\frac{\sum_{i\in S}(M-\frac{p_{i}}{s})}{nM-1}=\frac{1}{2},$
which is a contradiction. Thus, EqCardPartition must be a no instance.
Given a solution $S\subseteq[n]$ for $c$-FairnessScores, we can verify in polynomial time if $|S|=c$, compute SVD as well as check if $\sum_{i\in S}\alpha_{i}\geq\theta$ and $\sum_{i\in S}\beta_{i}\geq\theta$. Thus, $c$-FairnessScores is NP-complete. Naturally, $c$-FairnessScores reduces to MinFairnessScores, as the latter finds the smallest $c$ such that a solution to $c$-FairnessScores exists. Thus, MinFairnessScores is NP-hard. ∎
Even though MinFairnessScores is NP-hard, a $2$-factor approximation is trivial: sort ${\ell_{i}^{(k)}(A)}$’s in decreasing order and choose indices with highest leverage scores until they sum to $\theta$. Repeat the same for ${\ell_{i}^{(k)}(B)}$’s. This results in at most $2c$ columns, where $c$ is the optimal number of columns to satisfy $\theta$.
Input: $P=\{(\alpha_{1},\beta_{1}),\ldots,(\alpha_{n},\beta_{n})\}$, $\theta$
Output: $S\subseteq P$
1
[…]
3
 // add $(\alpha_{j},\beta_{j})$ to $S$ until either of the thresholds is satisfied
4 while $\sum_{(\alpha_{i},\beta_{i})\in S}\alpha_{i}<\theta$ and $\sum_{(\alpha_{i},\beta_{i})\in S}\beta_{i}<\theta$ do
5       $(\alpha_{j},\beta_{j})\leftarrow\max_{(\alpha_{j},\beta_{j})\in P\setminus S}(\alpha_{j}+\beta_{j})$
6       $S\leftarrow S\cup(\alpha_{j},\beta_{j})$
7
 // find $Q\in P\setminus S$ such that the other threshold is satisfied
8 if $\sum_{(\alpha_{i},\beta_{i})\in S}\alpha_{i}\geq\theta$ then
9       $Q\leftarrow\operatorname*{\arg\!\min}_{|Q|}\left(\sum_{j\in Q}\beta_{j}\geq\theta\right)$
10else
11       $Q\leftarrow\operatorname*{\arg\!\min}_{|Q|}\left(\sum_{j\in Q}\alpha_{j}\geq\theta\right)$
12
[…]
return $S$
Algorithm 1 FairScoresSampler
Next, we present an algorithm for MinFairnessScores that returns at most $\lceil 3c/2\rceil+1$ columns ($\approx 1.5$-approximation). [^1] An additional column ($+1$) is required in case $c$ is odd. The pseudocode is in Algorithm 1. For ease of notation, we denote $\alpha_{i}={\ell_{i}^{(k)}(A)}$ and $\beta_{i}={\ell_{i}^{(k)}(B)}$. The algorithm proceeds in two stages. In the first stage, at each iteration add to $S$ the index $i$ such that the cumulative gain $\alpha_{i}+\beta_{i}$ is maximized until a step $t\leq n$, where at least one of the inequalities is satisfied, i.e., either $\sum_{i\in S}\alpha_{i}\geq\theta$ or $\sum_{i\in S}\beta_{i}\geq\theta$. In the second step, sort the tuples of leverage scores in $[n]\setminus S$ based on their contribution to the unsatisfied inequality, in descending order. Finally, pick the rest of the tuples based on this order, until the threshold $\theta$ is satisfied. The following theorem establishes our approximation result. Note that the approximation is in terms of the number of columns $c$ in the optimal solution, and we have already established an $(1-\epsilon)^{-1/2}$ approximation in terms of our objective function.
Theorem 3.
Algorithm 1 returns a solution of at most $\lceil\frac{3}{2}c\rceil+1$ columns for MinFairnessScores, where $c$ is the number of columns in the optimal solution.
Proof.
Given $P=\{(\alpha_{1},\beta_{1}),\ldots,(\alpha_{n},\beta_{n})\}$, the task is to find the smallest subset $S\subseteq P$ such that $\sum_{i\in S}\alpha_{i}\geq\theta$ and $\sum_{i\in S}\beta_{i}\geq\theta$. At each iteration we select a tuple with the maximum contribution, i.e., $\alpha_{j}+\beta_{j}$ until some step $t$, where either $\sum_{i\in S}\alpha_{i}\geq\theta$ or $\sum_{i\in S}\beta_{i}\geq\theta$ is satisfied. Without loss of generality, at step $t$ we assume that $\sum_{(\alpha_{i},\beta_{i})\in S}\alpha_{i}\geq\theta$. Let $S^{*}$ be the optimal solution and $(\alpha_{i}^{*},\beta_{i}^{*})$ denote the contribution of $i$-th tuple in $S^{*}$. We assume $S^{*}$ is sorted in decreasing order according to $\alpha_{i}^{*}+\beta_{i}^{*}$. We note that this assumption makes our analysis easier without losing generality.
First we establish that $t\leq c$. Assume for contradiction that $t>c$ and that $\sum_{i=1}^{t}\alpha_{i}<\theta$ and that $\sum_{i=1}^{t}\beta_{i}<\theta$. Thus the algorithm has not terminated. From the optimality of $\alpha_{i}+\beta_{i}$ at step $t$ it holds that,
$\sum_{i=1}^{t}\alpha_{i}$ $+\sum_{i=1}^{t}\beta_{i}\geq\sum_{i=1}^{t}\alpha_{i}^{*}+\sum_{i=1}^{t}\beta_{i}^{*},$
$\sum_{i=1}^{t}\alpha_{i}$ $\geq\sum_{i=1}^{c}\alpha_{i}^{*}+\sum_{i=1}^{c}\beta_{i}^{*}-\sum_{i=1}^{t}\beta_{i}\geq 2\theta-\sum_{i=1}^{c}\beta_{i}\geq\theta,$
which is a contradiction, since we assumed $\sum_{i=1}^{t}\alpha_{i}<\theta$. Thus if $\sum_{i=1}^{t}\alpha_{i}\geq\theta$ then $1\leq t\leq c$. We now discern two cases for the value of $t$.
Case $t\leq\lceil\frac{c}{2}\rceil+1$: We have $\sum_{i=1}^{t}\alpha_{i}\geq\theta$. To satisfy the second inequality we choose tuples in $P\setminus S$ in decreasing order according to $\beta_{i}$, which is at most $c$ columns, since the optimal solution has $c$ columns. So, we have a solution with size $|S|\leq t+c=\lceil\frac{3}{2}c\rceil+1$.
Case $\lceil\frac{c}{2}\rceil+1<t\leq c$: Again, from the optimality of $\alpha_{i}+\beta_{i}$ at each step in $1\leq t\leq c$ we have,
$\sum_{i=1}^{t-1}\alpha_{i}+\sum_{i=1}^{t-1}\beta_{i}$ $\geq\sum_{i=1}^{t-1}\alpha_{i}^{*}+\sum_{i=1}^{t-1}\beta_{i}^{*}$
$\geq\sum_{i=1}^{\lceil\frac{c}{2}\rceil}\alpha_{i}^{*}+\sum_{i=\lceil\frac{c}{2}\rceil+1}^{t-1}\alpha_{i}^{*}+\sum_{i=1}^{\lceil\frac{c}{2}\rceil}\beta_{i}^{*}+\sum_{i=\lceil\frac{c}{2}\rceil+1}^{t-1}\beta_{i}^{*}$
$\geq\theta+\sum_{i=\lceil\frac{c}{2}\rceil+1}^{t-1}\alpha_{i}^{*}+\sum_{i=\lceil\frac{c}{2}\rceil+1}^{t-1}\beta_{i}^{*}$
The third step follows from the assumption that the optimal solution $S^{*}$ has tuples sorted in decreasing order according to $\alpha_{i}^{*}+\beta_{i}^{*}$. We also observe that $\sum_{i=1}^{t-1}\alpha_{i}=\theta-\alpha_{t}$. Therefore we have
$\sum_{i=1}^{t-1}\beta_{i}\geq\sum_{i=\frac{c}{2}+1}^{t-1}\alpha_{i}^{*}+\sum_{i=\frac{c}{2}+1}^{t-1}\beta_{i}^{*}+\alpha_{t}$
Suppose $t=\lceil\frac{c}{2}\rceil+1+q$. This means we can afford to add at most $c-q$ columns to our solution if we want to satisfy our bound on the cardinality of $S$.
At this stage, the algorithm picks the columns with the largest values of $\beta_{i}^{*}$. This means that from those in the optimal solution, we miss at most $q$ $\beta_{i}^{*}$’s from among the bottom ones. From above, we have
$\sum_{i=1}^{t-1}\beta_{i}\geq\sum_{i=\frac{c}{2}+1}^{t-1}\alpha_{i}^{*}+\sum_{i=\frac{c}{2}+1}^{t-1}\beta_{i}^{*}\geq\sum_{i=c-q}^{c}\alpha_{i}^{*}+\sum_{i=c-q}^{c}\beta_{i}^{*}\geq\sum_{i=c-q}^{c}\beta_{i}^{*}.$
This holds since tuples of the optimal solution are sorted in decreasing order of the value of the pair sums, which implies that the value we miss from not adding the last $q$ $\beta_{i}^{*}$’s is already covered by what we had, $\sum_{i=1}^{t-1}\beta_{i}$, so $\sum_{i=1}^{|S|}\beta_{i}\geq\sum_{i=1}^{t-1}\beta_{i}+\sum_{i=1}^{c-q}\beta_{i}^{*}\geq\theta.$ So the solution has at most $t-1+c-q=\lceil\frac{c}{2}\rceil+q+c-q=\lceil\frac{3}{2}\rceil c+1$ columns and satisfies the threshold. ∎
Even though Algorithm 1 offers an upper bound on the number of columns in the optimal solution $c$, depending on the task at hand, it may be undesirable to obtain more than $k$ columns. CSS with exactly $k$ columns (the definition of Problem 1) is a well-studied problem and a wide range of algorithms have been developed for it. These algorithms are typically based on QR-decomposition. Motivated by this, in the following section we propose two QR-decomposition-related algorithms for FairCSS.
Recall the impossibility results from Section 3: unless we pick $2k$ columns it may be impossible to achieve a relative-error approximation in terms of the rank-$k$ reconstruction error. Thus, the following algorithms are heuristics. They can be used either directly for FairCSS-MinMax or part of a two-stage approach, that we describe in detail in Section 6.
5 Fair QR decompositions
Numerous practical algorithms for CSS originate from numerical linear algebra, often relying on QR decomposition with column pivoting.
Definition 3 (QR decomposition with column pivoting).
Given a matrix $M\in\mathbb{R}^{m\times n}$ with $m\geq n$ and an integer $k\leq n$. Matrix $M$ can be expressed as the product of an orthonormal matrix $Q\in\mathbb{R}^{m\times m}$ and an upper triangular matrix $R\in\mathbb{R}^{n\times n}$. More precisely,
$M\Pi=QR=Q\begin{pmatrix}R_{11}&R_{12}\\ 0&R_{22}\end{pmatrix},$
where $R_{11}\in\mathbb{R}^{k\times k}$, $R_{12}\in\mathbb{R}^{k\times(n-k)}$, $R_{22}\in\mathbb{R}^{(n-k)\times(n-k)}$ and $\Pi\in\mathbb{R}^{n\times n}$ is a permutation matrix.
Column pivoting involves finding a permutation matrix $\Pi$ for a given matrix $M$, such that $\|R_{22}\|_{F}$ is minimised. When this comes with certain guarantees, it leads to a rank revealing QR decomposition (RRQR), which forms the basis for various algorithms with approximation guarantees in CSS. Specifically, if we denote the first $k$ columns of the permutation matrix $\Pi$ as $\Pi_{k}$, then choosing the column subset $C=M\Pi_{k}$ guarantees that $loss(M,C)$ is bounded i.e., $\|M-P_{C}M\|_{F}=\|R_{22}\|_{F}$ (Boutsidis et al., 2009). We tailor Low-RRQR (L-RRQR), originally introduced by Chan and Hansen (1994), to accommodate the two-group fair setting.
We present a brief description of L-RRQR. We start with the QR-decomposition of matrix $M$ to obtain matrices $Q$ and $R$. We initialise $R_{11}=0,R_{22}=R$ and build $R_{11}$ incrementally through column pivoting: at each iteration we permute a column of $R_{22}$ to the first position, through $\Pi$. Then, we compute the QR-decomposition again, drop the first row and column of the resulting $R$, and proceed recursively on it. The SVD serves in finding the pivot column: if $v\in\mathbb{R}^{n}$ is the right singular vector corresponding to the largest singular value, then successive permutations such that $|(\Pi^{T}v)|_{1}=\|v\|_{\infty}$ lead to a provably small $\|R_{22}\|_{F}$.
Fair pivoting. Note that L-RRQR may introduce unfairness, since we factorize the matrix $M$, without considering the error of the two groups, $A$ and $B$ separately. So the pivoted columns may benefit only one group. To address this, we adapt the pivoting strategy of L-RRQR to benefit the group that suffers the worse reconstruction error. Thus, we perform simultaneous RRQR on $A$ and $B$ and at step-$i$ obtain the corresponding $Q^{A}(i),R^{A}(i)$ and $Q^{B}(i),R^{B}(i)$. We inspect the spectra of both $R^{A}(i)$ and $R^{B}(i)$ and choose the pivot based on the following strategy: we choose the right singular vector $v$ of either $R^{A}(i)$ or $R^{B}(i)$ corresponding to $\max\{\sigma_{1}(R^{A}_{22}(i)),\sigma_{1}(R^{B}_{22}(i))\}$, and select $\Pi$ such that $|(\Pi^{T}v)|_{1}=\|v\|_{\infty}$. The algorithm pseudocode is available in Algorithm 2.
1
2Input: QR factorizations $A\Pi^{A}=Q^{A}R^{A}$, $B\Pi^{B}=Q^{B}R^{B}$, $k$
Output: permutation $\Pi_{k}$
1:  for i=1,…,k do
32:     $R_{22}^{A}\leftarrow R^{A}[i:,i:]$, $R_{22}^{B}\leftarrow R^{B}[i:,i:]$
3:     $v\leftarrow$ $\max\{\sigma_{1}(R^{A}_{22}(i)),\sigma_{1}(R^{B}_{22}(i))\}$
4:     Compute permutation $P$ such that $|(P^{T}v)_{1}|=\|P^{T}v\|_{\infty}$
5:     Compute QR fact. $R^{A}_{22}P=Q^{A}_{1}\tilde{R}^{A}_{22}$ and $R^{B}_{22}P=Q^{B}_{1}\tilde{R}^{B}_{22}$
6:     $\Pi\leftarrow\Pi\begin{pmatrix}I&0\\ 0&P\end{pmatrix}$
7:     $R^{A}\leftarrow\begin{pmatrix}R^{A}_{11}&{Q_{1}^{A}}^{T}R^{A}_{12}\\ 0&\tilde{R}^{A}_{22}\end{pmatrix}$ and $R^{B}\leftarrow\begin{pmatrix}R^{B}_{11}&{Q_{1}^{B}}^{T}R^{B}_{12}\\ 0&\tilde{R}^{B}_{22}\end{pmatrix}$
8:  end for
return $S$
Algorithm 2 Fair L-RRQR
6 Experiments
This section describes the experimental setup, datasets used, and presents the experimental evaluation results.
Experimental setup. Our implementation is written in python. We use numpy, scipy and scikit-learn for preprocessing, linear algebra operations as well as parallelization. Experiments are executed on a compute node with 32 cores and 256GB of RAM. Our implementations are available as open source (Matakos et al., 2023).
Datasets. We use juvenile recidivism data (recidivism) from Catalunya (Tolan et al., 2019) and medical expenditure survey data 2015 (meps(Cohen et al., 2009), as well as various datasets from the UCI-ML repository (Dua and Graff, 2017):“heart-cleveland" (heart), “adult" (adult), “german-credit" (german), “credit-card" (credit), “student performance" (student), “compas-recidivism" (compas), “communities" (communities). Data is processed by removing protected attributes, converting categorical variables to one-hot encoding and normalizing each column to unit $L^{2}$-norm. Group membership is based on Sex, except for communities where group membership is majority white or a non-white community. Dataset statistics are reported in Table 1.
Table 1: Dataset statistics. $m_{A}$ and $m_{B}$ are the number of instances in groups $A$ and $B$, respectively. $n$ is the number of columns. $\gamma$($A$), $\gamma$($B$) is rank of $A$, $B$.
Dataset $n$ $m_{A}$ $m_{B}$ $\gamma$($A$) $\gamma$($B)$
heart 14 201 96 13 13
german 63 690 310 49 47
credit 25 18 112 11 888 24 24
student 58 383 266 42 42
adult 109 21 790 10 771 98 98
compas 189 9 336 2 421 167 73
communities 104 1 685 309 101 101
recidivism 227 1 923 310 175 113
meps 1 247 18 414 17 013 1 217 1 200
Experimental evaluation. Our experiments are in threefold. First, we assess the efficacy of the proposed algorithms in addressing the problem of FairCSS-MinMax. This evaluation involves comparing the performance of the proposed algorithms, considering various experimental setups. Second, we evaluate the effectiveness of the FairCSS-MinMax objective in selecting column subsets that result in fair reconstruction errors. Specifically, we compare the reconstruction errors of each group in the optimal solutions obtained using the vanilla CSS objective versus the FairCSS-MinMax objective. Last, we investigate the price of fairness. This entails verifying potential trade-offs or costs associated to attain fairness according to FairCSS-MinMax.
6.1 Algorithms Evaluation
[…]
Picking $c\geq k$ columns. Recall that FairScoresSampler chooses columns based on threshold $\theta$, and it can choose $c\geq k$ columns. In the first experiment, we evaluate the algorithms performance with respect to a specific low-rank subspace of $A$ and $B$ over different values of $c$. Thus, we evaluate the performance according to MinMaxLoss while keeping $\|A-A_{k}\|_{F}$ and $\|B-B_{k}\|_{F}$ fixed. We perform this for six largest datasets for $k=20$ (for meps, $k=50$). Figure 1 shows the results.
Figure 1: MinMaxLoss for different values of $c$ and fixed target rank $k$
Note that Low QR and High QR can only sample at most $\min(rank(A),rank(B))$ columns. Low QR shows relatively good performance compared to Greedy, which has the best performance. In meps, Greedy and High QR, due to their higher complexity did not terminate within 24 hours. Note that in meps, FairScoresSampler performs worse than Random for $c<300$, which implies the rank $50$ leverage scores do not decay quickly for $c<300$, and thus are not very informative. In Supplementary B, we plot the rank-$k$ leverage scores for datasets for different values of $k$.
Two-stage sampling. Recall that, FairCSS-MinMax asks for exactly $k$ columns. To combine the efficiency of FairScoresSampler and effectiveness of algorithms such as Greedy, we introduce and evaluate a two-stage sampling approach that returns exactly $k$-columns. Similar ideas have been explored in CSS literature (Boutsidis et al., 2008). In the first stage, we run FairScoresSampler that takes as input a threshold $\theta$ and the rank-$k$ leverage scores of $A$ and $B$, and returns $c\geq k$ columns. In the second stage, we run Low QR, High QR or Greedy on the subset of columns returned in the first stage, to obtain a column subset of size $k$. We refer to these methods as S-Low QR, S-High QR and S-Greedy.
Table 2: Performance comparison of algorithms.
Dataset $c$ $k$ MinMax Loss
S-Low QR S-High QR S-Greedy Low QR Greedy Random
communities 89 10 1.27323 1.50763 1.17121 1.25939 1.16976 1.47485
94 23 1.32462 1.54204 1.2717 1.41701 1.23969 1.65518
98 51 1.40469 1.57761 1.42658 1.40934 1.39823 2.95985
compas 118 10 1.04747 1.2738 1.03356 1.05041 1.03057 1.19477
165 19 1.08617 1.34321 1.05688 1.08617 1.0537 1.27599
177 37 1.37174 1.41811 1.14835 1.47138 1.1291 1.67851
adult 70 10 1.02345 1.09485 1.02111 1.02345 1.01768 1.05641
96 22 1.03347 1.12764 1.0374 1.03347 - 1.0589
103 49 1.07796 1.19301 1.40252 1.08317 - 1.0994
german 53 10 1.08088 1.30176 1.08488 1.07711 1.07349 1.14205
54 15 1.1439 1.34599 1.11798 1.11871 1.11088 1.1966
54 24 1.20605 1.38489 1.192 1.20246 1.18624 1.36138
recidivism 134 10 1.02485 1.17864 1.02313 1.02236 1.01483 1.12757
174 24 1.05569 1.29805 1.04054 1.05567 1.03202 1.22332
212 57 1.31688 1.52031 1.16871 1.27495 1.13311 1.59933
student 45 10 1.10833 1.42094 1.10559 1.11333 1.10597 1.17856
46 14 1.14467 1.39265 1.14375 1.15592 1.14361 1.26605
47 21 1.18932 1.5756 1.17832 1.209 1.18771 1.56265
meps 428 10 1.14642 1.83209 - 1.05601 - 1.17759
382 32 1.20093 1.85843 1.777 1.11665 - 1.26749
338 100 1.33916 1.70488 2.48787 - - 1.47403
Table 2 reports the results for the seven largest datasets for various values of $k$. We set $\theta=k-\frac{1}{2}$ in all cases, except meps, where $\theta=\frac{3k}{4}$ to reduce the number of sampled columns. Thus in all datasets (except meps) the resulting number of columns in the first stage, is a $\sqrt{2}$-approximation to the optimal number of columns $c$. Column “$c$" in Table 2 indicates the number of columns returned in the sampling stage. Note that in some cases the number of columns is required to satisfy $\theta$ is significantly large. For each experiment, the best performing algorithm is highlighted in bold.
We observe that Greedy performs better in most cases, but does not terminate always within twenty-four hours. S-Greedy is faster than Greedy, due to the sampling stage, and the objective values are close. On the other hand, S-High QR does not perform well in practice, though in theory we expect it to perform better for $k$ closer to $n$. Finally, we demonstrate that (see meps) small $k$ does not mean fewer columns are sampled in the first stage, because the lower-rank leverage scores decay faster; thus more columns are required to satisfy threshold $\theta$. A visualization of the decay of leverage scores is reported in the Supplementary B.
Refer to caption Refer to caption Refer to caption
Figure 2: Comparison of reconstruction error of CSS and FairCSS for groups $A$ and $B$.
Refer to caption Refer to caption Refer to caption
Figure 3: Price of fairness.
6.2 Evaluation of the fair CSS objective
We examine the imbalance in reconstruction errors of the “vanilla" CSS objective of Problem 1 and FairCSS-MinMax objective for groups $A$ and $B$. We compute the optimal solution for each objective for various $k$ through exhaustive enumeration. ${\sf opt}(M),{\sf opt}(A)$ and ${\sf opt}(B)$ denote the reconstruction error of vanilla CSS for matrices $M,A$ and $B$, respectively. ${\sf minmax}(M)={\sf minmax}(A,B)$ denote the reconstruction error of matrix $M$. Further, ${\sf fair}(A)$ and ${\sf fair}(B)$ denote the reconstruction errors of FairCSS-MinMax for groups $A$ and $B$, respectively. Lastly, ${\sf fair}(M)$ denote the reconstruction error of $M$ corresponding to the optimal solution of FairCSS-MinMax. We note that brute-force enumeration is expensive, even after parallelization and extensive optimization, so we only report results for heart, student and german datasets.
Fairness of solutions. In Figure 2 we compare ${\sf opt}(A)$, ${\sf opt}(B)$, that is, the optimal solution of vanilla CSS, and the optimal solution of FairCSS-MinMax ${\sf fair}(A),{\sf fair}(B)$. In most instances, we observe that the reconstruction errors of groups are disproportionate in vanilla CSS. However, the degree of imbalance is not uniform across datasets. One source of imbalance could be vastly different group sizes. However, as observed in student, $|A|>|B|$, but ${\sf opt}(A)<{\sf opt}(B)$. This further supports the need for sophisticated approaches to fairness in CSS, beyond mere normalization.
Price of fairness. Next, we verify how the fairness objective influences the quality of the solution in terms of reconstruction error. In Figure 3, we report the optimal solution of vanilla CSS and FairCSS-MinMax to assess the trade-off between fairness and reconstruction error. This analysis quantifies the extent to which we sacrifice reconstruction error to achieve our fairness objective. In most cases, we observe no significant difference in reconstruction error between vanilla CSS and FairCSS-MinMax, that is, ${\sf opt}(M)$ and ${\sf fair}(M)$, even though the value of ${\sf minmax}(M)$ is significantly higher than ${\sf opt}(M)$. Note that, the observations cannot be generalized across all datasets, and we cannot conclusively claim that there is no trade-off between fairness and reconstruction error. For instance, in the case of the heart dataset, we observe a significant difference between values of ${\sf opt}(M)$ and ${\sf fair}(M)$ for $k=5$.
7 Conclusion
We introduced a novel CSS variant for a fair setting when the matrix rows are partitioned into two groups. Our goal is to minimize the reconstruction error for both groups via a min-max objective. We utilized leverage scores to present an approximation algorithm when the column count is relaxed, and presented rank revealing QR-factorisation-based algorithms when the column count is fixed. Extensive experiments on real-world data validated the effectiveness of our approach in improving fairness. As future work, a natural direction is to extend our results to more than two groups. Also, improving the approximation ratio of MinFairnessScores requires further investigation.
Acknowledgements
Suhas Thejaswi acknowledges support from the European Research Council (ERC) under the European Union’s Horizon $2020$ research and innovation program (grant agreement No. $945719$) and the European Unions’s SoBigData++ Transnational Access Scholarship. Antonis Matakos acknowledges support from the Academy of Finland through the grant "Model Management Systems: Machine learning meets Database Systems"- MLDB (32511).
Appendix A PSEUDOCODE of Fair H-RRQR
High rank revealing QR-factorisation (H-RRQR) is similar to low rank revealing QR-factorisation (L-RRQR), for details of L-RRQR see Section 5 of the main paper. In H-RRQR, we begin with $R_{11}=R,R_{22}=0$ and recursively build $R_{22}$ by moving columns to the back. The fair variant of H-RRQR, at step-$i$, chooses the right singular vector $v$ corresponding to $\min\{\sigma_{i}(R^{A}_{11}(i)),\sigma_{i}(R^{B}_{11}(i))\}$, where $\sigma_{i}$ is the bottom singular value. Then, we construct a permutation $\Pi_{i+1}$ such that $|(\Pi_{i+1}^{T}v)|_{i}=\|v\|_{\infty}$.
1
2Input: QR factorizations $A\Pi^{A}=Q^{A}R^{A}$, $B\Pi^{B}=Q^{B}R^{B}$, $k$
Output: permutation $\Pi_{k}$
1:  for i=n,…,n-k+1 do
2:     $R_{11}^{A}\leftarrow R^{A}[:i,:i]$, $R_{11}^{B}\leftarrow R^{B}[:i,:i]$
3:     $v\leftarrow$ $\min\{\sigma_{i}(R^{A}_{11}(i)),\sigma_{i}(R^{B}_{11}(i))\}$
4:     Compute permutation $P$ such that $|(P^{T}v)_{i}|=\|P^{T}v\|_{\infty}$
5:     Compute QR fact. $R^{A}_{11}P=Q^{A}_{1}\tilde{R}^{A}_{11}$ and $R^{B}_{11}P=Q^{B}_{1}\tilde{R}^{B}_{11}$
6:     $\Pi\leftarrow\Pi\begin{pmatrix}P&0\\ 0&I\end{pmatrix}$
7:     $R^{A}\leftarrow\begin{pmatrix}\tilde{R}^{A}_{11}&{Q_{1}^{A}}^{T}R^{A}_{12}\\ 0&R^{A}_{22}\end{pmatrix}$ and $R^{B}\leftarrow\begin{pmatrix}\tilde{R}^{B}_{11}&{Q_{1}^{B}}^{T}R^{B}_{12}\\ 0&R^{B}_{22}\end{pmatrix}$
8:  end for
return $S$
Algorithm 3 Fair H-RRQR
Appendix B Decay of Leverage Scores
We plot the leverage scores of groups $A$ and $B$ for the experiments in Table 2. The leverage scores are sorted separately for the two groups and plotted in decreasing order of their value.
Figure 4: Leverage scores of $A$ and $B$ for Table 2